# StrucSum: Graph-Structured Reasoning for Long Document Extractive Summarization with LLMs

### v1

@@ Abstract
Large language models (LLMs) have shown strong performance in zero-shot summarization, but often struggle to model document structure and identify salient information in long texts. In this work, we introduce StrucSum, a training-free prompting framework that enhances LLM reasoning through sentence-level graph structures. StrucSum injects structural signals into prompts via three targeted strategies: Neighbor-Aware Prompting (NAP) for local context, Centrality-Aware Prompting (CAP) for importance estimation, and Centrality-Guided Masking (CGM) for efficient input reduction. Experiments on ArXiv, PubMed, and Multi-News demonstrate that StrucSum consistently improves both summary quality and factual consistency over unsupervised baselines and vanilla prompting. Notably, on ArXiv, it boosts FactCC and SummaC by 19.2 and 9.7 points, indicating stronger alignment between summaries and source content. These findings suggest that structure-aware prompting is a simple yet effective approach for zero-shot extractive summarization with LLMs, without any training or task-specific tuning. Our source code is released at https://github.com/HaohanYuan01/StrucSum.
@@
Haohan Yuan1  Sukhwa Hong2  Haopeng Zhang1 1ALOHA Lab, University of Hawaii at Manoa 2University of Hawaii at Hilo {haohany, haopengz}@hawaii.edu sukhwa@hawaii.edu
@@ 1 Introduction
Extractive summarization aims to identify the most salient sentences from a document to produce concise and factually reliable summaries. This approach is especially valuable in high-stakes domains such as scientific publishing, biomedical research, legal analysis, and policy documentation, where accuracy and traceability are critical Zhang et al. (2023a); Ruan et al. (2022). As long-form documents grow in length and complexity, effective summarization requires not only identifying salient content at the sentence level but also reasoning over discourse structure and capturing long-range dependencies Zhong and Litman (2025); Koh et al. (2022), which remain challenging for many existing methods.
@@
To capture document-level structure, prior work has represented documents as sentence-level graphs based on semantic, positional, or rhetorical relationships Mihalcea and Tarau (2004); Jin et al. (2020); Wang et al. (2020). These graph-based approaches offer interpretable mechanisms for sentence selection via centrality and coverage Erkan and Radev (2004); Cui et al. (2020). However, most existing approaches require supervised training with transformer encoders Li et al. (2020); Zhang et al. (2022); Zhao et al. (2024), which limits their scalability and adaptability across domains.
@@
In the era of advanced large language models (LLMs), systems such as GPT-4 and LLaMA3 have shown strong zero-shot performance on general summarization tasks Yuan and Zhang (2024). Their ability to generalize across domains without task-specific training makes them appealing alternatives to fully supervised models Kojima et al. (2023). However, LLMs are trained on unstructured text and often struggle to recognize salient content or maintain factual consistency in long or multi-section documents due to their lack of structural awareness Zhang et al. (2024a). Although prompting strategies like in-context learning and chain-of-thought have enhanced LLM reasoning in other domains Wei et al. (2023), prompting for structural reasoning in long-document extractive summarization remains unexplored.
@@
To address this challenge, we propose StrucSum, a graph-structured prompting framework that enhances LLMs for long-document extractive summarization in a zero-shot setting. StrucSum builds a Text-Attributed Graph (TAG) for each document, where sentences are nodes connected by semantic or positional relationships. Instead of tuning model parameters, we inject structural signals into LLM inputs through three complementary strategies: (1) Neighbor-Aware Prompting (NAP) enriches each sentence with its immediate neighbors to support local reasoning; (2) Centrality-Aware Prompting (CAP) appends sentence importance scores derived from graph centrality to guide global selection; and (3) Centrality-Guided Masking (CGM) filters out low-salience content to reduce input length. These strategies jointly expose different aspects of document structure, driving LLMs to generate more faithful and context-aware summaries without additional training. We focus on extractive summarization as it provides a clearer testbed for structural reasoning, where identifying salient sentences depends more explicitly on discourse-level relationships Zhang et al. (2023a). While our method is designed for the extractive setting, it is model-agnostic and can be extended to abstractive summarization, which we leave for future work.
We evaluate StrucSum on two long-document summarization benchmarks (ArXiv and PubMed) and one multi-document benchmark (Multi-News). Experimental results show that our framework consistently outperforms both unsupervised methods and LLMs with standard prompting across ROUGE, BERTScore, and factual consistency metrics. Notably, on the ArXiv dataset, our NAP strategy improves FactCC scores by 19.2 points and SummaC scores by 9.7 points, showing significant gains in factual reliability. We summarize our main contributions as follows:
@@
  • We propose StrucSum, a structure-aware prompting framework for extractive summarization that operates in a fully zero-shot, training-free setting with LLMs.
  • We introduce three complementary prompting strategies: NAP, CAP, and CGM, that inject local and global structural signals into LLM inputs to support more accurate and faithful sentence selection.
@@
  • We conduct extensive experiments across three long-form summarization benchmarks. Results show that StrucSum consistently improves both summary quality and factual consistency across multiple metrics. Our findings suggest the importance of incorporating structural signals to enhance LLM reasoning in zero-shot extractive summarization.
@@ 2.1 Graph-based Summarization
Graph-based models have been widely adopted in extractive summarization for their ability to capture cross-sentence relationships within a document. Their sparse structure brings scalability and flexibility, making them particularly suitable for long-form inputs. Early approaches constructed sentence-level graphs using lexical or semantic similarity, enabling algorithms such as TextRank to identify salient content Mihalcea and Tarau (2004). Building on this foundation, later work has incorporated Graph Neural Networks (GNNs) to better model both local and global dependencies Chhabra et al. (2024). MuchSUM, for example, introduces a multi-channel graph convolutional network that encodes semantic, centrality, and positional features through a bipartite word-sentence graph, outperforming BERT-based baselines Mao et al. (2022). Heterogeneous GNNs have further been explored to jointly model word- and sentence-level interactions Wang et al. (2020), while citation graphs have been used to capture cross-document relatedness in scientific summarization Chen et al. (2022).
@@ 2.2 Summarization with LLMs
Large Language Models (LLMs) have significantly advanced the field of text summarization, demonstrating strong performance in both zero-shot and few-shot settings. Models such as ChatGPT, and LLaMA2 have achieved state-of-the-art results on various benchmarks, including CNN/DailyMail, XSum, and Multi-News Zhang et al. (2024a). These models excel in generating coherent and contextually relevant summaries without the need for extensive task-specific fine-tuning. However, challenges remain. LLMs often struggle with factual consistency, especially in abstractive summarization tasks, leading to potential hallucinations or misrepresentations of source content Pu et al. (2023). Additionally, their performance can vary across domains, with specialized areas such as medical or legal texts posing unique difficulties Yuan and Zhang (2024). In response, evaluation methods leveraging LLMs themselves Song et al. (2024), such as G-Eval, have been proposed to provide more nuanced assessments of summary quality Liu et al. (2023). Although recent work has begun exploring extractive summarization with LLMs in a training-free setting Zhang et al. (2023a), research in this area remains limited, particularly for long-form documents. Our work fills this gap by introducing structural signals to enhance the reasoning capabilities of LLMs, enabling more effective extractive summarization without any additional training or supervision.
@@ 3.1 Problem Definition
$\hat{\mathcal{S}}=f(P).\quad\text{where }\hat{\mathcal{S}}\subseteq\mathbf{S}$ (2)
@@ 3.3 Vanilla Prompting
Prompt Format.
The vanilla prompt is formatted as:
You are an expert in extractive summarization. Your task is to select the most important sentences from a document. On average, select around 4 key sentences. Sentence 1: "<Text_S1>" Sentence 2: "<Text_S2>" Sentence N: "<Text_SN>"
This setup defines the summarization task purely through textual input, representing the document as an ordered list $\mathbf{S}=(s_{1},s_{2},\dots,s_{N})$. This prompt template is reused and expanded upon in all structure-aware strategies. We use this formulation as the foundation for exploring how additional structural signals—introduced in later sections—can guide LLMs toward more accurate and faithful extractive summaries.
@@ 3.4 Structure-Aware Prompting
Sentence 1: "<Text_S1>" - 1-hop Neighbors: Sentence 2, 6, 7, 9, 13, 14, … Sentence 2: "<Text_S2>" - 1-hop Neighbors: Sentence 1, 4, 6, 9, 10, 15, …
By explicitly listing 1-hop neighbors, this format provides the LLM with local contextual cues that highlight semantically related sentences, helping it better assess each sentence’s role within its immediate discourse neighborhood.
@@
where $e_{ij}=1$ if an edge exists between $s_{i}$ and $s_{j}$, and 00 otherwise.
@@
Sentence 1 (Centrality: 0.67): "<Text_S1>" Sentence 2 (Centrality: 0.42): "<Text_S2>"
@@
Sentence 1: "<Text_S1>" Sentence 4: "<Text_S4>" Sentence 6: "<Text_S6>" [ Masked Sentences: 2, 3, 5, 7, 8, 9, 10, … ]
Input: Sentence set $\mathbf{S}=\{s_{1},\dots,s_{N}\}$;
Text-Attributed Graph $\mathcal{G}$;
Centrality coverage threshold $\rho\in(0,1)$
Output: Prompt $P$ with selected top-ranked sentences
1
$\triangleright$ Compute centrality $c_{i}$ for each $s_{i}\in\mathbf{S}$ using $\mathcal{G}$
2 for $i=1$ to $N$ do
3       $c_{i}\leftarrow\text{Centrality}(s_{i},\mathcal{G})$
4 end for
5
$\triangleright$ Select sentences covering $\rho$ of total centrality
6 Sort $\mathbf{S}$ by $c_{i}$ descending into $(s_{(1)},\dots,s_{(N)})$;
7 $C_{cumulative}\leftarrow 0$; $C_{target}\leftarrow\rho\cdot\sum_{k=1}^{N}c_{k}$;
8 for $j=1$ to $N$ do
9       $C_{cumulative}\leftarrow C_{cumulative}+c_{(j)}$;
10       if $C_{cumulative}\geq C_{target}$ then
11             break
12       end if
13
14 end for
15
$\triangleright$ Build prompt $P$ with top $j$ sentences
16 Initialize prompt $P\leftarrow[\;]$;
17 for $k=1$ to $j$ do
18       Add text of $s_{(k)}$ to $P$;
19
20 end for
21
return $P$
Algorithm 1 Centrality-Guided Masking
The full procedure is detailed in Algorithm 1. This approach offers a more efficient encoding of structural relevance by reducing input length while preserving key informational content. Full prompt examples for vanilla prompting, NAP, CAP, and CGM are provided in Appendix A for clarity and reproducibility.
@@ 4.1 Experimental Setup
Lead 8.69 26.51 0.8310 13.26 30.69 0.8388 8.83 25.79 0.8481
LexRank (2004) 6.04 22.32 0.8094 10.46 28.69 0.8313 7.15 23.21 0.8364
TextRank (2004) 8.69 23.37 0.8163 12.20 29.27 0.8329 9.32 25.57 0.8435
PacSum (2019) 9.66 26.01 0.8197 12.61 31.45 0.8446 7.89 24.62 0.8408
@@
We experiment with three LLMs: GPT4o-mini and GPT4o from OpenAI, and LLaMA3.1-70B-Instruct from Meta. GPT models are accessed via the official OpenAI API [^1] https://openai.com/index/openai-api/, while LLaMA3 is accessed through the NVIDIA API [^2] https://build.nvidia.com/models. All models are used in zero-shot settings.
@@
We evaluate summary quality using a combination of standard, faithfulness-oriented, and holistic metrics. For content overlap, we report ROUGE-2 and ROUGE-L Lin (2004), which capture n-gram overlap and longest common subsequence between system and reference summaries. For semantic similarity, we report BERTScore F1 Zhang et al. (2020), which measures contextual embedding similarity between summaries and references. To assess factual consistency, we use FactCC Kryściński et al. (2019), a classifier-based metric that detects factual errors in generated summaries, and SummaC Laban et al. (2022), which estimates sentence-level consistency using natural language inference models. Finally, we report G-Eval Liu et al. (2023), a GPT-based evaluation framework that provides human-aligned scores across multiple dimensions, including coherence, consistency, fluency, and relevance. G-Eval offers a more comprehensive view of overall summary quality beyond traditional overlap-based metrics.
@@
To construct the TAG, we first encode each sentence using the pre-trained all-MiniLM-L6-v2 Sentence-BERT model. We then compute cosine similarity between all sentence pairs and add an undirected edge if their similarity exceeds a tunable threshold $\theta$. All hyperparameters are tuned on the development set. We report the values of three hyperparameters used in our framework. The settings are summarized in Table 3.
@@ 4.2 Summary Quality Evaluation
Table 2 reports the performance of StrucSum and baseline methods on three long-document summarization datasets: ArXiv, PubMed, and Multi-News. Results are grouped into unsupervised extractive methods and LLM-based approaches, with the latter further divided into vanilla prompting and structure-aware prompting. Across all datasets, structure-aware prompting consistently improves over vanilla LLM prompting, highlighting the value of incorporating document structure.
Among unsupervised baselines, Lead performs best on PubMed and remains competitive on ArXiv and Multi-News. Graph-based methods like LexRank, TextRank, and PacSum offer varying degrees of improvement, but all are outperformed by LLM-based methods—particularly on scientific texts where structural cues are more critical.
For LLMs, Neighbor-Aware Prompting (NAP) yields the most consistent and substantial improvements across datasets and metrics. Centrality-Aware Prompting (CAP) also brings notable gains, while Centrality-Guided Masking (CGM) shows moderate improvements with a favorable trade-off in input length. Across models, GPT-4o-mini achieves the strongest overall results when combined with structural prompting. These findings confirm that incorporating structural signals enhances the extractive summarization capability of LLMs, especially for long-form documents.

### v2

@@ Abstract
Large language models (LLMs) have shown strong performance in zero-shot summarization, but often struggle to model document structure and identify salient information in long texts. In this work, we introduce StrucSum, a training-free prompting framework that enhances LLM reasoning through sentence-level graph structures. StrucSum injects structural signals into prompts via three targeted strategies: Neighbor-Aware Prompting (NAP) for local context, Centrality-Aware Prompting (CAP) for importance estimation, and Centrality-Guided Masking (CGM) for efficient input reduction. Experiments on ArXiv, PubMed, and Multi-News demonstrate that StrucSum consistently improves both summary quality and factual consistency over unsupervised baselines and vanilla prompting. In particular, on ArXiv, it increases FactCC and SummaC by 19.2% and 8.0% points, demonstrating stronger alignment between summaries and source content. The ablation study shows that the combination of multiple strategies does not yield clear performance gains; therefore, structure-aware prompting with graph-based information represents a promising and underexplored direction for the advancement of zero-shot extractive summarization with LLMs. Our source code is publicly available.  [^1] https://github.com/HaohanYuan01/StrucSum
@@
Haohan Yuan1   Sukhwa Hong2   Haopeng Zhang1 thanks: corresponding author 1ALOHA Lab, University of Hawaii at Manoa 2University of Hawaii at Hilo {haohany, sukhwa, haopengz}@hawaii.edu
@@ 1 Introduction
Despite recent advances in LLM-based extractive summarization Zhang et al. (2023a), summarizing long and structured documents remains challenging for large language models due to limited context windows and insufficient discourse awareness Zhang et al. (2024a). As documents increase in length and complexity, effective summarization requires not only identifying salient sentences but also reasoning over discourse structure and capturing long-range dependencies Zhong and Litman (2025); Koh et al. (2022), which continue to pose difficulties for existing approaches.
@@
Extractive summarization selects key sentences to create concise, factually reliable summaries, crucial for domains requiring accuracy and traceability such as science, biomedicine, law, and policy. Zhang et al. (2023a); Ruan et al. (2022). To capture document-level structure, prior work has explored representing documents as sentence trees or graphs based on semantic, positional, or rhetorical relationships Mihalcea and Tarau (2004); Jin et al. (2020); Wang et al. (2020), offering interpretable mechanisms for sentence selection via centrality and coverage Erkan and Radev (2004); Cui et al. (2020). However, most existing approaches require supervised training with human written gold summaries Li et al. (2020); Zhang et al. (2022); Zhao et al. (2024), which limits their scalability and adaptability across domains.
@@
Recently, LLMs have shown strong zero-shot performance in general summarization tasks Yuan and Zhang (2024). Their ability to generalize across domains without task-specific fine-tuning makes them attractive alternatives to fully supervised models Kojima et al. (2023). However, LLMs are trained on unstructured text and often struggle to recognize salient content and maintain factual consistency in long-form documents due to their lack of structural awareness Zhang et al. (2024a). Although prompting strategies like in-context learning Brown et al. (2020) and chain-of-thought Wei et al. (2023) have enhanced LLM reasoning in other domains, how to incorporate structural reasoning in prompting for long-document summarization remains unexplored.
@@
In this paper, we propose StrucSum, a structure-aware prompting method for long-document extractive summarization in a fully zero-shot setting with LLMs. StrucSum constructs a Text-Attributed Graph (TAG) where sentences are nodes linked by semantic relations, and introduces three strategies to inject structural signals into model inputs: (1) Neighbor-Aware Prompting (NAP) adds local sentence context; (2) Centrality-Aware Prompting (CAP) appends graph-based importance scores; and (3) Centrality-Guided Masking (CGM) removes low-salience content to reduce length. These strategies enhance structural reasoning without any parameter updates. While developed for extractive summarization, StrucSum is model-agnostic and can be readily extended to abstractive or multi-document settings. Across datasets and LLMs, StrucSum improves both summary quality and factuality, revealing that symbolic graph priors can guide black-box LLMs without training. Our main contributions include:
@@
  • We propose StrucSum, a structure-aware prompting framework for extractive summarization in a zero-shot, training-free LLM setting, featuring three structural strategies: NAP, CAP, and CGM.
@@
  • Experiments on three long-form summarization benchmarks show that StrucSum consistently improves both summary quality and factual consistency over standard prompting.
  • We find that different structural strategies support different goals: CGM is most effective for maximizing overlap with reference summaries, NAP best maintains factual consistency, and CAP offers a balance between performance and input length. These findings provide guidance for selecting strategies based on task needs.
@@ 2.1 Graph-based Summarization
Graph-based models are a fundamental approach in extractive summarization, widely used for capturing cross-sentence relationships. Early work, such as TextRank Mihalcea and Tarau (2004), constructed sentence graphs based on lexical similarity to rank and select salient content. Subsequent research integrated Graph Neural Networks (GNNs) to learn more expressive, data-driven representations of these document structures Chhabra et al. (2024). This has led to the development of more complex graph designs for capturing finer-grained semantics, including the bipartite word-sentence graphs in MuchSUM Mao et al. (2022) and the joint modeling of words and sentences in heterogeneous GNNs Wang et al. (2020). These techniques have also been adapted for specific contexts: HGSUM Li et al. (2023) uses a compressed heterogeneous graph for multi-document summarization, G-SEEK Moro et al. (2023) employs a lightweight selector for low-resource settings, and GSS constructs citation graphs for scientific articles Chen et al. (2022).
@@ 2.2 Summarization with LLMs
Large Language Models have shown strong zero-shot performance on summarization benchmarks like CNN/DailyMail and XSum Zhang et al. (2024a). Despite these advances, a main challenge is factual consistency, as LLMs are prone to generating hallucinations unfaithful to the source content Pu et al. (2023). Furthermore, their effectiveness often decreases in specialized domains, such as medical or legal texts Yuan and Zhang (2024). To mitigate these issues, QA-prompting Sinha (2025) has been introduced to address positional biases by employing question-answering to extract key information before final summary generation. To address domain-specific difficulties, Qi et al. (2025) improves summarization for medical texts by combining prompt optimization and supervised fine-tuning of LLMs. In addition, evaluation frameworks like G-Eval Liu et al. (2023); Song et al. (2024) further enable finer-grained assessment of summary quality. Although some efforts explore training-free extractive summarization with LLMs Zhang et al. (2023a), this area remains underdeveloped, particularly for long documents. Our work contributes to improving LLM reasoning in extractive summarization through zero-shot learning using structural signals, thereby eliminating the need for additional training or supervision.
@@ 3.1 Problem Definition
[…]
@@ 3.3 Vanilla Prompting
Prompt Format
The vanilla prompt defines the task and represents the document as an ordered list $\mathbf{S}=(s_{1},s_{2},\dots,s_{N})$. This prompt template is reused and expanded upon in all structure-aware strategies.
System Instruction: You are an expert in extractive summarization. Your task is to select the most important sentences from a document. Guideline: On average, select $k$ key sentences. Sentence List: Sentence 1: “The UN Security Council met to discuss …” Sentence 2: “Several nations expressed concern over …” Sentence 3: “The meeting concluded with a call for …” $\dots$ Sentence N: “International observers were invited to …” Expected Output Format: { "selected_sentences": [1, 3, 5] }
@@ 3.4 Structure-Aware Prompting
Context: Each sentence is followed by its 1-hop neighbors. Sentence 1: “<Text_S1>” Neighbors: Sentence 2, 6, 7, 9, 13, 14 Sentence 2: “<Text_S2>” Neighbors: Sentence 1, 4, 6, 9, 10, 15 $\dots$
By explicitly listing 1-hop neighbors, this format provides the LLM with local contextual cues that highlight semantically related sentences, helping it better assess the role of each sentence within its immediate discourse neighborhood.
@@
where $e_{ij}=1$ if an edge exists between $s_{i}$ and $s_{j}$, and 0 otherwise.
@@
Context: Each sentence is presented with its centrality score. Sentence 1 (Centrality: 0.82): "<Text_S1>" Sentence 2 (Centrality: 0.51): "<Text_S2>" $\dots$
@@
Context: Only top-ranked sentences (by centrality) are shown in full; others are masked. Sentence 1: "<Text_S1>" Sentence 4: "<Text_S4>" Sentence 6: "<Text_S6>" $\dots$
The full procedure is detailed in Algorithm 1. This approach offers a more efficient encoding of structural relevance by reducing input length while preserving key informational content. Full prompt examples for vanilla prompting, NAP, CAP, and CGM are provided in the supplementary materials for clarity and reproducibility.
Input: Sentences $\mathbf{S}=\{s_{1},\dots,s_{N}\}$; Graph $\mathcal{G}$; Threshold $\rho\in(0,1)$
Output: Prompt $P$ with top-ranked sentences
1 Compute centrality score $c_{i}$ for each $s_{i}$ using $\mathcal{G}$;
2 Sort $\mathbf{S}$ by $c_{i}$ in descending order;
3 Select minimal prefix $\{s_{(1)},\dots,s_{(j)}\}$ such that $\sum_{k=1}^{j}c_{(k)}\geq\rho\cdot\sum_{i=1}^{N}c_{i}$;
4 Build prompt $P$ with full text of selected sentences $\{s_{(1)},\dots,s_{(j)}\}$;
5 Mask the remaining sentences;
6 return $P$;
Algorithm 1 Centrality-Guided Masking (CGM)
@@ 4.1 Experimental Setup
Lead 8.69 26.51 83.10 13.26 30.69 83.88 8.83 25.79 84.81
LexRank (2004) 6.04 22.32 80.94 10.46 28.69 83.13 7.15 23.21 83.64
TextRank (2004) 8.69 23.37 81.63 12.20 29.27 83.29 9.32 25.57 84.35
PacSum (2019) 9.66 26.01 81.97 12.61 31.45 84.46 7.89 24.62 84.08
@@
We experiment with three LLMs: GPT4o-mini and GPT4o from OpenAI, and LLaMA3.1-70B-Instruct from Meta. GPT models are accessed via the official OpenAI API [^2] https://openai.com/index/openai-api/, while LLaMA3 is accessed through the NVIDIA API [^3] https://build.nvidia.com/models. For calling LLMs, we consistently set the temperature to 0 and used the default “top_p” value of $1.0$. The “max_tokens” parameter was set to $100$ for all models to control the output length.
@@
We evaluate summary quality using a combination of standard, faithfulness-oriented, and holistic metrics. For content overlap, we report ROUGE-2 and ROUGE-L Lin (2004), which capture n-gram overlap and longest common subsequence between system and reference summaries. For semantic similarity, we report BERTScore F1 Zhang et al. (2020), which measures contextual embedding similarity between summaries and references. To assess factual consistency, we use FactCC kryściński2019evaluatingfactualconsistencyabstractive, a classifier-based metric that detects factual errors in generated summaries, and SummaC Laban et al. (2022), which estimates sentence-level consistency using natural language inference models. Finally, we report G-Eval Liu et al. (2023), a GPT-based evaluation framework that provides human-aligned scores across multiple dimensions, including coherence, consistency, fluency, and relevance. G-Eval offers a more comprehensive view of overall summary quality beyond traditional overlap-based metrics.
@@
To construct the TAG, we first encode each sentence using the pre-trained all-MiniLM-L6-v2 Sentence-BERT model. We then compute cosine similarity between all sentence pairs and add an undirected edge if their similarity exceeds a tunable threshold $\theta$. All hyperparameters are tuned on the development set. The values used for each dataset are provided in Appendix 4.2.
4.2 Hyperparameter Settings
We report the values of three hyperparameters used in our framework. The settings are summarized in Table 3.
@@ 4.3 Summary Quality Evaluation
Table 2 reports the performance of StrucSum and baseline methods on three datasets: ArXiv, PubMed, and Multi-News. We compare unsupervised extractive baselines with LLM-based methods, including both vanilla prompting and structure-aware prompting. We observe that, in most cases, structure-aware prompting outperforms vanilla LLM prompting.
Among the baselines, Lead performs best on PubMed (R-2: 13.26) and remains competitive elsewhere. Graph-based methods like LexRank and PacSum show varying results but fall short of LLM performance, particularly on long documents.
For LLMs, all three structural strategies improve over vanilla prompting. On Multi-News, GPT4o-mini with NAP achieves an R-2 score of 10.13, representing the largest relative gain (+0.67) over its vanilla baseline. On ArXiv, CAP gives the highest R-2 score (15.39) and BERTScore (84.77) among GPT4o-mini variants. On PubMed, CGM performs best, with an R-2 of 18.21 and BERTScore of 84.92.
Overall, structure-aware prompting improves summary quality across different datasets. NAP provides the largest relative improvements, while CAP and CGM achieve the highest absolute scores on specific datasets and metrics.